OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. y = 1/2 x + 4 y = 1/2 x y = 1/2 x + 2

OpenStudy (dtan5457):

slope intercept = y=mx+b all you need to do is solve for b, the y intercept

OpenStudy (dtan5457):

plug the values in so you get 2=4(1/2)+b

Nnesha (nnesha):

you would get same answer but it will be better to write x in parenthesis instead of slope y = mx+b ==> y =1/2(x)+b :)

OpenStudy (anonymous):

so would it be y = 1/2 x + 2

OpenStudy (anonymous):

????

OpenStudy (dtan5457):

I AM HERE

OpenStudy (dtan5457):

is 1/2x+2 your answer?

OpenStudy (anonymous):

yeh

OpenStudy (dtan5457):

2=4(1/2)+b 2=2+b what does b equal..?

OpenStudy (dtan5457):

1/2x+2 is not right

OpenStudy (dtan5457):

b=0, you must know the equation now lol